Determine the amount of heat needed to be added to (2.70x10^0) moles of diatomic oxygen in an isobaric process to change its temperature from T_i=(9.900x10^1) K to T_f=(4.2500x10^2) K.

Provide your answer in joules, in scientific notation, use 3 significant figures in your answer. The value of the specific heat at constant pressure for diatomic oxygen is 29.2

Note: Your answer is assumed to be reduced to the highest power possible.

The answer I provided was:

257x10^2 J

I got this question wrong. Please explain why this is wrong, how to get the correct answer and what the correct answer is.

Thank you!
Answer #1

given
n = 2.7 mol
Ti = 99 K
Tf = 425 K
Cp = 29.2 J/(mol K)

The amount of heat needed,

Q = n*Cp*delta_T

= 2.7*29.2*(425 - 99)

= 25702 J

= 2.57*10^4 J   <<<<<<<<---------------Answer

Another method
Workdone by the gas in Isobaric process,

W = P*(Vf - Vi)

= n*R*(Tf - Ti)

= 2.70*8.314*(425 - 99)

= 7318 J

change in internal energy, delta_U = (5/2)*n*R*(Tf - Ti)

= (5/2)*2.7*8.314*(425 - 99)

= 18295 J

now use first law of thermodynamics, Q = W + delta_U

= 7318 + 18295

= 25613 J

= 2.56*10^4 J <<<<<<<<---------------Answer